ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

9. Ordinary Differential Equations and Dynamical Systems

1991 MS Classification: 58 Kopanskii, Alexander, Institute of Mathematics, Kishinev, Moldova Normal forms of vector fields satisfying certain geometric conditions The talk describes the simplified polynomial resonant normal forms of vector fields satisfying some geometric conditions (namely, preserving volume or symplectic forms or contact structures) with respect to C^{k} smooth changes of variables preserving the same structure.